# From Efficiency to Adaptivity: A Deeper Look at Adaptive Reasoning in Large Language Models

### v1

@@ Abstract
Mingchen Gao University at Buffalo mgao8@buffalo.edu          Zhenyi Wang University of Central Florida zhenyi.wang@ucf.efu
@@ 1 Introduction
To realize adaptive reasoning, LLMs must flexibly employ different reasoning strategies depending on task demands. Reasoning in LLMs encompasses diverse cognitive patterns, among which three classical paradigms—inductive, deductive, and abductive reasoning—provide a useful conceptual lens. Inductive reasoning generalizes from observed examples, as seen in in-context learning brown2020language; deductive reasoning applies explicit logical rules to derive valid conclusions, underlying structured chain-of-thought wei2022chain and program-of-thought chen2022program frameworks; abductive reasoning seeks plausible explanations for observed outcomes. However, rather than organizing methods by reasoning type, this survey focuses on how adaptivity is implemented in practice. We propose a systematic taxonomy distinguishing training-based approaches, which internalize adaptive allocation strategies through reinforcement learning, supervised fine-tuning, and learned controllers, from training-free approaches, which achieve adaptivity through prompt conditioning, feedback-driven halting, and modular composition at inference time. Our contributions are threefold:
@@ 2.1 What is reasoning in LLMs?
Reasoning is a cornerstone of human intelligence, enabling inference and decision-making from evidence and rules. Following huang-chang-2023-towards and 10.1145/3729218, reasoning can be viewed as a logical and systematic process that uses prior knowledge and contextual evidence to arrive at valid conclusions or hypotheses. Within the context of LLMs, reasoning refers to the model’s capacity to generate and verify intermediate representations that connect an input prompt $\mathbf{x}$ to an output $y$ through a structured inference procedure.
@@ 2.1.1 Reasoning as Conditional Generation
Following bandyopadhyay2025thinkingmachinessurveyllm, A language model defines a conditional probability distribution over output sequences:
@@
Extending this formulation, phan2023training formalize reasoning as a latent-variable conditional generation process. Specifically, the model can be viewed as first generating a latent reasoning trace $\mathbf{r}$, which is an unobserved sequence of intermediate thought step, and then producing the final output $\mathbf{y}$ conditioned on both $\mathbf{x}$ and $\mathbf{r}$. This leads to the marginal likelihood:
@@ 2.1.2 Reasoning as Resource-Bounded Optimization
Following alomrani2025reasoningbudgetsurveyadaptive, test-time reasoning can be viewed as a trade-off between computation and performance. Two equivalent optimization views are commonly considered: a constrained form ($L_{1}$) that fixes the compute budget, and a penalized form ($L_{2}$) that adaptively balances accuracy and efficiency. We adopt the $L_{2}$ formulation, which naturally aligns with the idea of adaptive reasoning:
@@ 2.1.3 Limitations of Efficient Reasoning
LLMs often exhibit an overthinking phenomenon, producing unnecessarily long chains of thought (CoT) for simple problems, thereby increasing compute without improving accuracy (Sui2025StopOverthinking; alomrani2025reasoningbudgetsurveyadaptive). Efficiency-oriented methods attempt to limit reasoning tokens, steps, or decoding paths to reduce test-time compute. However, such static control allocates identical budgets to all inputs, regardless of difficulty or confidence, leading to redundant reasoning on easy cases and insufficient reasoning on complex ones.
Empirical studies confirm that fixed-length CoT models often waste computation on easy inputs while failing on hard ones (Sui2025StopOverthinking). The key limitation is that computation is treated as an external constraint rather than a self-regulated process: intelligent reasoning should dynamically decide how much to think based on uncertainty and task complexity.
@@ 2.2 Forms of Reasoning
A reasoning process generates an intermediate trajectory $\mathbf{r}=(r_{1},\ldots,r_{K})$ that mediates between input $\mathbf{x}$ and output $\mathbf{y}$, where each step $r_{k}$ is autoregressively conditioned on $(\mathbf{x},r_{<k})$. We call a reasoning episode $R_{\bm{\theta}}(\mathbf{x})=(\mathbf{r},\mathbf{y})$ valid if it satisfies (1) Causal coherence — each step $r_{k}$ depends only on $\mathbf{x}$ and previous steps, ensuring a forward information flow; (2) Goal consistency — the reasoning trajectory $\mathbf{r}$ produces an answer $\mathbf{y}$ that faithfully and correctly fulfills the task objective, following the principles of faithful and step-wise reasoning (lee2025evaluating; lyu2023faithful)
@@
This ensures that the reasoning trajectory preserves logical validity at every step (ling2023deductive). In LLMs, deductive reasoning is approximated through verifiable chain-of-thought generation, where each step can be checked for logical entailment from previous statements (ling2023deductive; SealsShalin2024Deductive; Xia2025FormalLogic).
@@
This formalizes Occam’s Razor (Baker2022-BAKSSE-2) that preferring hypotheses that explain data with minimal complexity. When $C(h)=-\log p(h)$, this reduces to Maximum A Posteriori (MAP) estimation under Bayesian induction. In LLMs, inductive reasoning manifests primarily in in-context learning (ICL) (brown2020language), where models extrapolate implicit rules from few-shot demonstrations to unseen cases (Ye2025ICLAttribution). However, empirical studies show that LLMs do not always follow simplicity principles in practice (SunSaparov2025Occam).
@@
This instantiates inference to the best explanation (Baker2022-BAKSSE-2): among hypotheses that entail observation $O$ when combined with background knowledge $\mathcal{B}$, abduction selects the most probable one. In LLMs, abductive reasoning is realized through conditional generation where the model samples from $p_{\theta}(H\mid O)$ to produce plausible explanations. Benchmarks such as Abductive (bhagavatula2020abductivecommonsensereasoning) and INABHYD (SunSaparov2025Occam) evaluate this ability by measuring explanation coherence and plausibility.
@@ 2.3 Adaptive Reasoning
where $\lambda$ controls the cost penalty and $\mathcal{D}$ is the data distribution. The optimization over $\bm{\phi}\in\Phi$ captures two complementary approaches to adaptive reasoning. In training-based methods3.1), $\bm{\phi}$ is learned during training, for instance, IBPO (Yu2025IBPO) trains a neural network to predict reasoning budgets, while C3oT (kang2025c3ot) learns to generate short or long reasoning traces conditionally. In training-free methods3.2), $\bm{\phi}$ represents fixed inference-time control such as entropy-based halting (laaouach2025halt) or adaptive sampling thresholds, applied without parameter updates. This unified formulation clarifies how adaptivity arises from either learned policies or inference-time heuristics, both of which modulate $\pi_{\bm{\theta}}$ to allocate reasoning effort based on input characteristics.
Unlike static efficiency methods that apply uniform compute budgets to all inputs, this framework enables input-dependent resource allocation by conditioning reasoning on $\bm{\phi}(\mathbf{x})$, which adapts to task difficulty and uncertainty. Recent work further shows that such adaptivity can emerge even without reinforcement learning: karan2025reasoning demonstrates that inference-time adaptive sampling alone can elicit strong reasoning behaviors from base LLMs, highlighting that reasoning adaptivity may arise naturally from sampling dynamics rather than explicit policy optimization.
@@
(1) Training-free adaptivity. Adaptivity can emerge without retraining the model, by dynamically regulating reasoning length or sampling during inference. laaouach2025halt halt generation when the stepwise entropy $H_{i}=-\!\sum_{a}p_{i}(a)\log p_{i}(a)$ drops below a threshold $\theta$. aggarwal2023letssamplestepstep use a Dirichlet prior over samples to detect when majority confidence exceeds $C_{\text{thresh}}$, stopping further decoding. Han2025TALE estimates minimal token budgets per instance and predicts per-sample cost as $\text{Cost}=\frac{1}{N}\sum_{i=1}^{N}T(M(x_{i}))$. These methods achieve instance-level compute control through entropy, confidence, or budget signals.
(2) Training-based adaptivity. Other approaches learn a reasoning policy that internalizes computation allocation during optimization. Yu2025IBPO formulate inference-aware optimization with explicit budget constraints, where the policy is trained to maximize reward while satisfying correctness thresholds that implicitly controls compute allocation. Such training-based strategies generalize TTC to the learning phase, yielding policies that adapt depth and effort before deployment.
@@ 3 Taxonomy of Adaptive Reasoning
Training-based3.1)Training-free3.2) RL Methods (§3.1.1): IBPO Yu2025IBPO, LCPO Aggarwal2025LCPO, Reasoning-efficient RL Arora2025EfficientRL, SABER Zhao2025SABER, LSPO Chen2025LSPO, LASER liu2025learnreasonefficientlyadaptive, SFPO Li2025SFPO, BRPO qi2025optimizinganytimereasoningbudget, HBPO lyu2025hierarchicalbudgetpolicyoptimization,BudgetThinker wen2025budgetthinker,AdaReasoner Tan2025AdaReasoner SFT & Distillation Methods (§3.1.2): C3oT kang2025c3ot, CoT-Valve Ma2025CoTValve, TokenSkip Xia2025TokenSkip, LS-Mixture SFT Yu2025LSMixtureSFT, Distilling 2-1 Zhang2024DistillingTwoOne, SoftCoT Xu2025SoftCoT, SoftCoT++ Xu2025SoftCoTpp, Skip-Thinking chen2025skipthinkingchunkwisechainofthoughtdistillation, MoR xiong2025mixturereasoningsteachlarge, LIFT-COT electronics14081662, TALE Han2025TALE, SCoT wang2025efficientreasoningllmsspeculative Learned Controllers and Routers Methods (§3.1.3): RouteLLM ong2024routellm, SelfBudgeter Li2025SelfBudgeter, ThinkSwitcher Liang2025ThinkSwitcher,AdaMOE  zeng2024adamoe,Route to Reason (pan2025routereasonadaptiverouting) Prompt-conditioned Methods (§3.2.1): Token complexity Lee2025TokenComplexity, Constrained-CoT Nayab2025ConciseThoughts, Sketch-of-Thought aytes2025sketchofthoughtefficientllmreasoning, Chain-of-Draft Xu2025ChainOfDraft, ConciseHint Tang2025ConciseHint, Plan-and-Budget Lin2025PlanBudget Feedback-driven Methods (§3.2.2): Entropy halting laaouach2025halt, DEER Yang2025DEER, ES-CoT mao2025earlystoppingchainofthoughtslarge, EAGER Scalena2025EAGER, Self-Certainty Kang2025SelfCertainty, Sequence entropy zhu2025uncertainty; sharma2025think, DeepConf Fu2025DeepConf, Answer convergence liu2025answerconvergence, Filler suppression wang2025nowait, Speculative Search wang2025specreasoning, Value-Guided Search wang2025valueguided, Adaptive-Consistency aggarwal2023letssamplestepstep, RASC wan2025reasoningawareselfconsistencyleveraging, Fast BoN Sun2024FastBoN, ST-BoN Wang2025STBoN, DPTS Ding2025DPTS, Sample Smart Li2025SampleSmartNotHard,MetaReasoner sui2025metareasonerdynamicguidanceoptimized, Modular Methods (§3.2.3): AlphaOne Zhang2025AlphaOne, Bag-of-Tricks liu2025bagtricksinferencetimecomputation, L2S merging Wu2025L2S, CoThink fan2025cothink,ACM Yao2025ACM, Lang-specific merging Pipatanakul2025LangSpecificMerge, Kimi k1.5 TeamKimi2025Kimi15
@@ 3.1.1 Reinforcement Learning Methods
Reinforcement-learning (RL)–based approaches integrate reasoning accuracy and computational efficiency into a unified optimization objective. IBPO (Yu2025IBPO) formulates an inference-budget–constrained policy optimization problem where the policy learns to allocate larger token budgets to harder questions and smaller ones to easier instances. LCPO (Aggarwal2025LCPO) directly adds a length-adherence(i.e., a reward that penalizes deviations from the target reasoning length) reward term to policy gradients, enabling the agent to match target reasoning lengths and revealing the emergence of short-reasoning models that preserve logical fidelity. (Arora2025EfficientRL) propose a reasoning-efficient RL framework that shapes rewards to favor concise yet correct reasoning trajectories. SABER (Zhao2025SABER) introduces switchable and balanced reward schedules to train models under multiple budget tiers. LSPO (Chen2025LSPO) develops a length-aware data sampling strategy that dynamically filters training batches based on response length (e.g., retaining the shortest and longest) to stabilize training. Other variants include LASER (liu2025learnreasonefficientlyadaptive), which introduces adaptive, difficulty-aware reward shaping, penalizing lengthy reasoning more heavily on simple tasks, and Slow-Fast Policy Optimization (SFPO) (Li2025SFPO) structures each update into fast, reposition, and slow phases to reduce gradient noise and stabilize convergence. Multi-objective approaches such as BRPO (qi2025optimizinganytimereasoningbudget) optimizes anytime reasoning via budget-sampled dense rewards and a variance-reduced policy gradient, and HBPO (lyu2025hierarchicalbudgetpolicyoptimization) structures exploration across hierarchical token-budget groups to preserve capability while reducing average tokens. Together, these works demonstrate how reinforcement learning can encode budget awareness and trade-offs directly into learned reasoning policies through principled reward shaping and constraint design. Other variants include BudgetThinker (wen2025budgetthinker) introduces dynamic control tokens whose embeddings encode remaining compute, enabling the model to adjust reasoning depth continuously during decoding, while uniquely combining an RL-based reward function with explicit control tokens, bridging the gap between purely implicit policy optimization and supervised control. AdaReasoner (Tan2025AdaReasoner) learns an adaptive policy via reinforcement learning to dynamically configure reasoning parameters (e.g., temperature, step length, prompt type) for different tasks

### v2

# arXiv:2511.10788v2[cs.AI] 01 Mar 2026

@@ Abstract
Mingchen Gao University at Buffalo mgao8@buffalo.edu          Yu Tian University of Central Florida yu.tian2@ucf.edu          Zhenyi Wang University of Central Florida zhenyi.wang@ucf.efu
@@ 1 Introduction
To realize adaptive reasoning, LLMs must flexibly employ different reasoning strategies depending on task demands. Reasoning in LLMs encompasses diverse cognitive patterns, among which three classical paradigms—inductive, deductive, and abductive reasoning—provide a useful conceptual lens. Inductive reasoning generalizes from observed examples, as seen in in-context learning Brown et al. (2020); deductive reasoning applies explicit logical rules to derive valid conclusions, underlying structured chain-of-thought Wei et al. (2022) and program-of-thought Chen et al. (2022) frameworks; abductive reasoning seeks plausible explanations for observed outcomes. However, rather than organizing methods by reasoning type, this survey focuses on how adaptivity is implemented in practice. We propose a systematic taxonomy distinguishing training-based approaches, which internalize adaptive allocation strategies through reinforcement learning, supervised fine-tuning, and learned controllers, from training-free approaches, which achieve adaptivity through prompt conditioning, feedback-driven halting, and modular composition at inference time. Our contributions are threefold:
@@ 2.1 What is reasoning in LLMs?
Reasoning is a cornerstone of human intelligence, enabling inference and decision-making from evidence and rules. Following Huang and Chang (2023) and Sun et al. (2025), reasoning can be viewed as a logical and systematic process that uses prior knowledge and contextual evidence to arrive at valid conclusions or hypotheses. Within the context of LLMs, reasoning refers to the model’s capacity to generate and verify intermediate representations that connect an input prompt $\mathbf{x}$ to an output $y$ through a structured inference procedure.
@@ 2.1.1 Reasoning as Conditional Generation
Following Bandyopadhyay et al. (2025), A language model defines a conditional probability distribution over output sequences:
@@
Extending this formulation, Phan et al. (2023) formalize reasoning as a latent-variable conditional generation process. Specifically, the model can be viewed as first generating a latent reasoning trace $\mathbf{r}$, which is an unobserved sequence of intermediate thought step, and then producing the final output $\mathbf{y}$ conditioned on both $\mathbf{x}$ and $\mathbf{r}$. This leads to the marginal likelihood:
@@ 2.1.2 Reasoning as Resource-Bounded Optimization
Following Alomrani et al. (2025), test-time reasoning can be viewed as a trade-off between computation and performance. Two equivalent optimization views are commonly considered: a constrained form ($L_{1}$) that fixes the compute budget, and a penalized form ($L_{2}$) that adaptively balances accuracy and efficiency. We adopt the $L_{2}$ formulation, which naturally aligns with the idea of adaptive reasoning:
@@ 2.1.3 Limitations of Efficient Reasoning
LLMs often exhibit an overthinking phenomenon, producing unnecessarily long chains of thought (CoT) for simple problems, thereby increasing compute without improving accuracy (Sui et al., 2025a; Alomrani et al., 2025). Efficiency-oriented methods attempt to limit reasoning tokens, steps, or decoding paths to reduce test-time compute. However, such static control allocates identical budgets to all inputs, regardless of difficulty or confidence, leading to redundant reasoning on easy cases and insufficient reasoning on complex ones.
Empirical studies confirm that fixed-length CoT models often waste computation on easy inputs while failing on hard ones (Sui et al., 2025a). The key limitation is that computation is treated as an external constraint rather than a self-regulated process: intelligent reasoning should dynamically decide how much to think based on uncertainty and task complexity.
@@ 2.2 Forms of Reasoning
A reasoning process generates an intermediate trajectory $\mathbf{r}=(r_{1},\ldots,r_{K})$ that mediates between input $\mathbf{x}$ and output $\mathbf{y}$, where each step $r_{k}$ is autoregressively conditioned on $(\mathbf{x},r_{<k})$. We call a reasoning episode $R_{\bm{\theta}}(\mathbf{x})=(\mathbf{r},\mathbf{y})$ valid if it satisfies (1) Causal coherence — each step $r_{k}$ depends only on $\mathbf{x}$ and previous steps, ensuring a forward information flow; (2) Goal consistency — the reasoning trajectory $\mathbf{r}$ produces an answer $\mathbf{y}$ that faithfully and correctly fulfills the task objective, following the principles of faithful and step-wise reasoning (Lee and Hockenmaier, 2025; Lyu et al., 2023)
@@
This ensures that the reasoning trajectory preserves logical validity at every step (Ling et al., 2023). In LLMs, deductive reasoning is approximated through verifiable chain-of-thought generation, where each step can be checked for logical entailment from previous statements (Ling et al., 2023; Seals and Shalin, 2024; Xia et al., 2025b).
@@
This formalizes Occam’s Razor (Baker, 2022) that preferring hypotheses that explain data with minimal complexity. When $C(h)=-\log p(h)$, this reduces to Maximum A Posteriori (MAP) estimation under Bayesian induction. In LLMs, inductive reasoning manifests primarily in in-context learning (ICL) (Brown et al., 2020), where models extrapolate implicit rules from few-shot demonstrations to unseen cases (Ye et al., 2025). However, empirical studies show that LLMs do not always follow simplicity principles in practice (Sun and Saparov, 2025).
@@
This instantiates inference to the best explanation (Baker, 2022): among hypotheses that entail observation $O$ when combined with background knowledge $\mathcal{B}$, abduction selects the most probable one. In LLMs, abductive reasoning is realized through conditional generation where the model samples from $p_{\theta}(H\mid O)$ to produce plausible explanations. Benchmarks such as Abductive (Bhagavatula et al., 2020) and INABHYD (Sun and Saparov, 2025) evaluate this ability by measuring explanation coherence and plausibility.
@@ 2.3 Adaptive Reasoning
where $\lambda$ controls the cost penalty and $\mathcal{D}$ is the data distribution. The optimization over $\bm{\phi}\in\Phi$ captures two complementary approaches to adaptive reasoning. In training-based methods3.1), $\bm{\phi}$ is learned during training, for instance, IBPO (Yu et al., 2025b) trains a neural network to predict reasoning budgets, while C3oT (Kang et al., 2025a) learns to generate short or long reasoning traces conditionally. In training-free methods3.2), $\bm{\phi}$ represents fixed inference-time control such as entropy-based halting (Laaouach, 2025) or adaptive sampling thresholds, applied without parameter updates. This unified formulation clarifies how adaptivity arises from either learned policies or inference-time heuristics, both of which modulate $\pi_{\bm{\theta}}$ to allocate reasoning effort based on input characteristics.
Unlike static efficiency methods that apply uniform compute budgets to all inputs, this framework enables input-dependent resource allocation by conditioning reasoning on $\bm{\phi}(\mathbf{x})$, which adapts to task difficulty and uncertainty. Recent work further shows that such adaptivity can emerge even without reinforcement learning: Karan and Du (2025) demonstrates that inference-time adaptive sampling alone can elicit strong reasoning behaviors from base LLMs, highlighting that reasoning adaptivity may arise naturally from sampling dynamics rather than explicit policy optimization.
@@
(1) Training-free adaptivity. Adaptivity can emerge without retraining the model, by dynamically regulating reasoning length or sampling during inference. Laaouach (2025) halt generation when the stepwise entropy $H_{i}=-\!\sum_{a}p_{i}(a)\log p_{i}(a)$ drops below a threshold $\theta$. Aggarwal et al. (2023) use a Dirichlet prior over samples to detect when majority confidence exceeds $C_{\text{thresh}}$, stopping further decoding. Han et al. (2025) estimates minimal token budgets per instance and predicts per-sample cost as $\text{Cost}=\frac{1}{N}\sum_{i=1}^{N}T(M(x_{i}))$. These methods achieve instance-level compute control through entropy, confidence, or budget signals.
(2) Training-based adaptivity. Other approaches learn a reasoning policy that internalizes computation allocation during optimization. Yu et al. (2025b) formulate inference-aware optimization with explicit budget constraints, where the policy is trained to maximize reward while satisfying correctness thresholds that implicitly controls compute allocation. Such training-based strategies generalize TTC to the learning phase, yielding policies that adapt depth and effort before deployment.
@@ 3 Taxonomy of Adaptive Reasoning
Training-based3.1)Training-free3.2) RL Methods (§3.1.1): IBPO Yu et al. (2025b), LCPO Aggarwal and Welleck (2025), Reasoning-efficient RL Arora and Zanette (2025), SABER Zhao et al. (2025), LSPO Chen et al. (2025a), LASER Liu et al. (2025b), SFPO Wang et al. (2025g), BRPO Qi et al. (2025), HBPO Lyu et al. (2025),BudgetThinker Wen et al. (2025),AdaReasoner Wang et al. (2025d) SFT & Distillation Methods (§3.1.2): C3oT Kang et al. (2025a), CoT-Valve Ma et al. (2025), TokenSkip Xia et al. (2025a), LS-Mixture SFT Yu et al. (2025a), Distilling 2-1 Yu et al. (2024), SoftCoT Xu et al. (2025b), SoftCoT++ Xu et al. (2025c), Skip-Thinking Chen et al. (2025b), MoR Xiong et al. (2025), LIFT-COT Chen and Li (2025), TALE Han et al. (2025), SCoT Wang et al. (2025b) Learned Controllers and Routers Methods (§3.1.3): RouteLLM Ong et al. (2025), SelfBudgeter Li et al. (2025b), ThinkSwitcher Liang et al. (2025),AdaMOE  Zeng et al. (2024),Route to Reason (Pan et al., 2025) Prompt-conditioned Methods (§3.2.1): Token complexity Lee et al. (2025), Constrained-CoT Nayab et al. (2025), Sketch-of-Thought Aytes et al. (2025), Chain-of-Draft Xu et al. (2025a), ConciseHint Tang et al. (2025), Plan-and-Budget Lin et al. (2025) Feedback-driven Methods (§3.2.2): Entropy halting Laaouach (2025), DEER Yang et al. (2025), ES-CoT Mao et al. (2025), EAGER Scalena et al. (2025), Self-Certainty Kang et al. (2025b), Sequence entropy Zhu et al. (2025); Sharma and Chopra (2025), DeepConf Fu et al. (2025), Answer convergence Liu and Wang (2025), Filler suppression Wang et al. (2025a), Speculative Search Wang et al. (2025f), Value-Guided Search Wang et al. (2025c), Adaptive-Consistency Aggarwal et al. (2023), RASC Wan et al. (2025), Fast BoN Sun et al. (2024), ST-BoN Wang et al. (2025e), DPTS Ding et al. (2025), Sample Smart Li et al. (2025a),MetaReasoner Sui et al. (2025b), Modular Methods (§3.2.3): AlphaOne Zhang et al. (2025), Bag-of-Tricks Liu et al. (2025a), L2S merging Wu et al. (2025), CoThink Fan et al. (2025),ACM Yao et al. (2025), Lang-specific merging Pipatanakul et al. (2025), Kimi k1.5 Kimi (2025)
@@ 3.1.1 Reinforcement Learning Methods
Reinforcement-learning (RL)–based approaches integrate reasoning accuracy and computational efficiency into a unified optimization objective. IBPO (Yu et al., 2025b) formulates an inference-budget–constrained policy optimization problem where the policy learns to allocate larger token budgets to harder questions and smaller ones to easier instances. LCPO (Aggarwal and Welleck, 2025) directly adds a length-adherence(i.e., a reward that penalizes deviations from the target reasoning length) reward term to policy gradients, enabling the agent to match target reasoning lengths and revealing the emergence of short-reasoning models that preserve logical fidelity. (Arora and Zanette, 2025) propose a reasoning-efficient RL framework that shapes rewards to favor concise yet correct reasoning trajectories. SABER (Zhao et al., 2025) introduces switchable and balanced reward schedules to train models under multiple budget tiers. LSPO (Chen et al., 2025a) develops a length-aware data sampling strategy that dynamically filters training batches based on response length (e.g., retaining the shortest and longest) to stabilize training. Other variants include LASER (Liu et al., 2025b), which introduces adaptive, difficulty-aware reward shaping, penalizing lengthy reasoning more heavily on simple tasks, and Slow-Fast Policy Optimization (SFPO) (Wang et al., 2025g) structures each update into fast, reposition, and slow phases to reduce gradient noise and stabilize convergence. Multi-objective approaches such as BRPO (Qi et al., 2025) optimizes anytime reasoning via budget-sampled dense rewards and a variance-reduced policy gradient, and HBPO (Lyu et al., 2025) structures exploration across hierarchical token-budget groups to preserve capability while reducing average tokens. Together, these works demonstrate how reinforcement learning can encode budget awareness and trade-offs directly into learned reasoning policies through principled reward shaping and constraint design. Other variants include BudgetThinker (Wen et al., 2025) introduces dynamic control tokens whose embeddings encode remaining compute, enabling the model to adjust reasoning depth continuously during decoding, while uniquely combining an RL-based reward function with explicit control tokens, bridging the gap between purely implicit policy optimization and supervised control. AdaReasoner (Wang et al., 2025d) learns an adaptive policy via reinforcement learning to dynamically configure reasoning parameters (e.g., temperature, step length, prompt type) for different tasks